# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v2

@@ 5.2 Large scale geometry of thin manifolds
The previous result fails if we remove the lower bounds on the Ricci curvature in conditions (38) and (40), thanks to the gluing results in [GromovLawsonSurgery1, GromovLawsonSurgery2, WolfsonKricci]. Indeed, consider the gluing of four copies of $S^{2}\times[0,+\infty)$ to $S^{3}$ minus four disjoint small balls. By [WolfsonKricci], we can equip this space with a Riemannian metric $g$ with the following properties:
@@
[…]
At the same time, Theorem 5.6 fails on $M$. With a similar construction using [GromovLawsonSurgery1, GromovLawsonSurgery2], one shows that also the Ricci bound in (40) cannot be removed.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $C,\delta>0$ satisfying the following. Let $(M^{n},g,p)$ be a manifold satisfying one of the conditions (38), (39), or (40). Then
@@
Proof of 1. For $t\in(0,1]$, the claim is a consequence of the classical volume comparison results for manifolds with Ricci curvature bounded from below by $-1$. Hence, we only consider the case $t>1$.
Let $x\in X$ and let $\phi(s_{x})$ be a projection of $x$ on $\phi(I)$. Let $\{s_{i}\}_{i=1}^{\lceil t\rceil}$ be a $1$-net of $B^{I}_{t}(s_{x})$. By triangular inequality, $B_{t}(x)$ is contained in the $10C^{\prime}$-neighbourhood of $\phi(B^{I}_{t}(s_{x}))$. Hence, again by triangle inequality and our choice of $\{s_{i}\}_{i=1}^{\lceil t\rceil}$, it follows
@@
By the aforementioned volume bounds for manifolds with Ricci curvature bounded from below by $-1$, there exists a constant $C>0$ (depending only on the dimension), such that $\mathsf{Vol}(B_{t}(x))\leq C\lceil t\rceil\leq 2Ct$.
@@
So we only need to consider the case when $M$ is non-compact and it is contained in the $C^{\prime}$ neighbourhood of a ray. Fix $x\in M$ and let $t_{x}\geq 0$ be such that $x\in B_{C^{\prime}}(r(t_{x}))$. By Brunn-Minkowski inequality [S2],
@@
To conclude the proof, we are left to show that $A_{1/2}(2t_{x})$ is contained in $B_{10C^{\prime}}(r(t_{x}))$. To this aim, consider a segment $l$ connecting points $a\in B_{1}(r(0))$ and $b\in B_{1}(r(2t_{x}))$. Let $t_{l}\geq 0$ be such that $r(t_{l})$ is a projection of $l(\mathrm{L}(l)/2)$ on $r$. Since $X$ is in the $C^{\prime}$-neighbourhood of $r$, by triangle inequality, it holds
@@
The previous result, in particular, shows that the manifolds in question, when $\mathsf{Ric}_{M}\geq 0$, fit into the framework of [ZhuLinear].
@@
An analog of Theorem 5.8 is not to be expected in higher dimension. Indeed a manifold with $\mathsf{R}_{3}\geq 1$ and almost non-negative Ricci curvature might have exponential volume growth at infinity (consider rescalings of products of the hyperbolic plane and a sufficiently small $2$-sphere). Similarly, there exists a manifold with $\mathsf{R}_{5}\geq 1$ and non-negative Ricci curvature with unit balls of arbitrarily small volume (consider a product of a $2$-sphere and the manifold constructed in [SormaniExample, Example $26$]).
Let us recall some properties of covering spaces. Let $\pi:\tilde{X}\to X$ be a covering between metric spaces. Let $p\in X$ and $\tilde{p}\in\pi^{-1}(p)$. Given a loop $\gamma$ in $p$, we can consider its lift $\tilde{\gamma}$ on $\tilde{X}$ with starting point $\tilde{p}$. This lift is unique, and homotopic loops give rise to homotopic lifts (see [Hatcher, p. 60]).
In particular, one can consider the endpoint of $\tilde{\gamma}$, which we will denote $M(\tilde{p},\gamma)$. This way of obtaining new points, given an element $[\gamma]\in\pi_{1}(X)$ and a point $\tilde{p}\in\tilde{X}$, is called the monodromy action of $\gamma$ on $\tilde{p}$. Since homotopic loops give rise to the same monodromy action, if $[\gamma]$ is the identity in $\pi_{1}(X)$, then $M(\tilde{p},\gamma)=\tilde{p}$.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $C,\delta>0$ satisfying the following. Let $(M^{n},g,p)$ be a manifold satisfying one of the conditions (38), (39), or (40). The following hold.
@@
    If there exists a loop $\gamma\subset M$ such that $[\gamma]\in\pi_{1}(M)$ has infinite order, then $M$ is contained in a $C$-neighbourhood of $\gamma$.
@@
Proof of 5. Let $\gamma\subset M$ be a loop such that $[\gamma]\in\pi_{1}(M)$ has infinite period. Let $\tilde{\gamma}$ be a lift of $\gamma$ to $\tilde{M}$, and let $\tilde{\gamma}^{\mathbb{N}}:[0,+\infty)\to\tilde{M}$ be a lift of the infinite curve $\sum_{i\in\mathbb{N}}\gamma$. Since $[\gamma]\in\pi_{1}(M)$ has infinite period, it holds
@@
Let $r_{\tilde{M}}$ be the ray or the line in $\tilde{M}$ whose $C^{\prime}$-neighbourhood contains $\tilde{M}$ (which exists combining Theorem 5.6 and Proposition 5.4). By (47) and Lemma 4.5, there exist $t_{1},t_{2}$ in the domain of $r_{\tilde{M}}$ with $t_{2}\geq 100C^{\prime}+100\mathrm{L}(\gamma)+t_{1}$ such that $\pi_{r_{\tilde{M}}}(\tilde{\gamma}^{\mathbb{N}})$ contains a $3C^{\prime}$-net of $[r_{\tilde{M}}(t_{1}),r_{\tilde{M}}(t_{2})]$. Assume now that there is a point $p\in M$ such that $\mathsf{d}(p,\gamma)\geq 10C^{\prime}$. Let $l$ be a segment realizing the distance from $p$ to $\gamma$ and let $p_{l}$ be the footpoint of $l$ on $\gamma$. The set of preimages of $p_{l}$ in $\tilde{M}$ via the covering map form a $\mathrm{L}(\gamma)$-net of $\tilde{\gamma}^{\mathbb{N}}$ in $\tilde{M}$. In particular, there is a preimage $\tilde{p}_{l}$ which is $10C^{\prime}+\mathrm{L}(\gamma)$ close to $r_{\tilde{M}}((t_{1}+t_{2})/2)$. Let $\tilde{l}$ be the lift of $l$ to $\tilde{M}$ starting from $\tilde{p}_{l}$ (we are orienting $l$ from $p_{l}$ to $p$). Then,
@@
At the same time, since $\tilde{p}_{l}$ is $10C^{\prime}+\mathrm{L}(\gamma)$ close to $r_{\tilde{M}}((t_{1}+t_{2})/2)$, it has a projection $r_{\tilde{M}}(t)$ on $r_{\tilde{M}}$ for $t\in[t_{1},t_{2}]$. Using that $r_{\tilde{M}}(t)$ is $7C^{\prime}$-close to a point in $\tilde{\gamma}^{\mathbb{N}}$ and (48), it holds
@@
The appendix shows how to obtain a result by Jiang-Naber [Naberconjectures, Theorem 2.17]. Since the result was claimed in [Naberconjectures] without proof, for the reader’s convenience, we provide an argument consisting in a combination of Theorem 1.1 with [CheegerNaber, Theorem 1.3]. The result is not used anywhere else in the paper, and it is included as a further application of the techniques developed in the manuscript.
@@
The next theorem corresponds to [CheegerNaber, Theorem 1.3].
@@
One can now deduce Jiang-Naber’s result [Naberconjectures, Theorem 2.17].

### v3

@@ 5.2 Large scale geometry of thin manifolds
The previous result fails if we remove the lower bounds on the Ricci curvature in conditions (30) and (32), thanks to the gluing results in [29, 28, 60]. Indeed, consider the gluing of four copies of $S^{2}\times[0,+\infty)$ to $S^{3}$ minus four disjoint small balls. By [60], we can equip this space with a Riemannian metric $g$ with the following properties:
@@
$\fint_{B_{1}(x)}\mathsf{R}_{2}\,d\mathsf{Vol}\geq\epsilon\quad\text{for every }x\in M.$
At the same time, Theorem 5.6 fails on $M$. With a similar construction using [29, 28], one shows that also the Ricci bound in (32) cannot be removed.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $C,\delta>0$ satisfying the following. Let $(M^{n},g)$ be a manifold satisfying one of the conditions (30), (31), or (32). Then
@@
Proof of 1. For $t\in(0,1]$, the claim is a consequence of the classical volume comparison results for manifolds with Ricci curvature bounded from below by 1-1. Hence, we only consider the case $t>1$.
Let $x\in X$ and let $\phi(s_{x})\in\pi_{\phi(I)}(x)$. Let $\{s_{i}\}_{i=1}^{\lceil t\rceil}$ be a $1$-net of $B^{I}_{t}(s_{x})$. By triangular inequality, $B_{t}(x)$ is contained in the $10C^{\prime}$-neighbourhood of $\phi(B^{I}_{t}(s_{x}))$. Hence, again by triangle inequality and our choice of $\{s_{i}\}_{i=1}^{\lceil t\rceil}$, it follows
@@
By the aforementioned volume bounds for manifolds with Ricci curvature bounded from below by 1-1, there exists a constant $C>0$ (depending only on the dimension), such that $\mathsf{Vol}(B_{t}(x))\leq C\lceil t\rceil\leq 2Ct$.
@@
So we only need to consider the case when $M$ is non-compact and it is contained in the $C^{\prime}$ neighbourhood of a ray. Fix $x\in M$ and let $t_{x}\geq 0$ be such that $x\in B_{C^{\prime}}(r(t_{x}))$. By Brunn-Minkowski inequality [53],
@@
To conclude the proof, we are left to show that $A_{1/2}(2t_{x})$ is contained in $B_{10C^{\prime}}(r(t_{x}))$. To this aim, consider a segment $l$ connecting points $a\in B_{1}(r(0))$ and $b\in B_{1}(r(2t_{x}))$. Let $t_{l}\geq 0$ be such that $r(t_{l})\in\pi_{r}\big(l(\mathrm{L}(l)/2)\big)$ on $r$. Since $X$ is in the $C^{\prime}$-neighbourhood of $r$, by triangle inequality, it holds
@@
The previous result, in particular, shows that the manifolds in question, when $\mathsf{Ric}_{M}\geq 0$, fit into the framework of [63].
@@
An analog of Theorem 5.8 is not to be expected in higher dimension. Indeed a manifold with $\mathsf{R}_{3}\geq 1$ and almost non-negative Ricci curvature might have exponential volume growth at infinity (consider rescalings of products of the hyperbolic plane and a sufficiently small $2$-sphere). Similarly, there exists a manifold with $\mathsf{R}_{5}\geq 1$ and non-negative Ricci curvature with unit balls of arbitrarily small volume (consider a product of a $2$-sphere and the manifold constructed in [52, Example $26$]).
Let us recall some properties of covering spaces. Let $\pi:\tilde{X}\to X$ be a covering between metric spaces. Let $p\in X$ and $\tilde{p}\in\pi^{-1}(p)$. Given a loop $\gamma$ in $p$, we can consider its lift $\tilde{\gamma}$ on $\tilde{X}$ with starting point $\tilde{p}$. This lift is unique, and homotopic loops give rise to homotopic lifts (see [33, p. 60]).
In particular, one can consider the endpoint of $\tilde{\gamma}$, which we will denote $M(\tilde{p},\gamma)$. This way of obtaining new points, given an element $\gamma]\in\pi_{1}(X)$ and a point $\tilde{p}\in\tilde{X}$, is called the monodromy action of $\gamma$ on $\tilde{p}$. Since homotopic loops give rise to the same monodromy action, if $\gamma]$ is the identity in $\pi_{1}(X)$, then $M(\tilde{p},\gamma)=\tilde{p}$.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $C,\delta>0$ satisfying the following. Let $(M^{n},g)$ be a manifold satisfying one of the conditions (30), (31), or (32). The following hold.
@@
    If there exists a loop $\gamma\subset M$ such that $\gamma]\in\pi_{1}(M)$ has infinite order, then $M$ is contained in a $C$-neighbourhood of $\gamma$.
@@
Proof of 5. Let $\gamma\subset M$ be a loop such that $\gamma]\in\pi_{1}(M)$ has infinite period. Let $\tilde{\gamma}$ be a lift of $\gamma$ to $\tilde{M}$, and let $\tilde{\gamma}^{\mathbb{N}}:[0,+\infty)\to\tilde{M}$ be a lift of the infinite curve $\sum_{i\in\mathbb{N}}\gamma$. Since $\gamma]\in\pi_{1}(M)$ has infinite period, it holds
@@
[…]
@@
At the same time, since $\tilde{p}_{l}$ is $10C^{\prime}+\mathrm{L}(\gamma)$ close to $r_{\tilde{M}}((t_{1}+t_{2})/2)$, it holds $r_{\tilde{M}}(t)\in\pi_{r_{\tilde{M}}}\big(\tilde{p}_{l}\big)$ for some $t\in[t_{1},t_{2}]$. Using that $r_{\tilde{M}}(t)$ is $7C^{\prime}$-close to a point in $\tilde{\gamma}^{\mathbb{N}}$ and (40), it holds
@@
The appendix shows how to obtain a result by Jiang-Naber [45, Theorem 2.17]. Since the result was claimed in [45] without proof, for the reader’s convenience, we provide an argument consisting in a combination of Theorem 1.1 with [15, Theorem 1.3]. The result is not used anywhere else in the paper, and it is included as a further application of the techniques developed in the manuscript.
@@
The next theorem corresponds to [15, Theorem 1.3].
@@
One can now deduce Jiang-Naber’s result [45, Theorem 2.17].